An ice cube is placed o top of an overturned spher
Answer #1

Here, N + mv2/R = mgcos(theta)

=>   applying consevation of energy

=>   1/2 * m * v2 = mgR * (1 - cos(theta))

=> putting N = 0

=>   v2/R = gcos(theta)

=> 2gR * (1 - cos(theta))/R = gcos(theta)

=>   2 * (1 - cos(theta)) = cos(theta)

=>    coa(theta) = 2/3

=> theta = 48.189 degree